A 200 g puck revolves in a circle on a frictionless table at the end of a 50.0 cm long string. The puck's angular momentum is 3.00 kg⋅m2/s.

Part A

What is the tension in the string?
Answer #1

Here,

m = 0.20 Kg

L = 0.50 m

angular momentum = 3 Kg.m^2/s

angular momentum , 3 = m * v * L

3 = 0.20 * 0.50 * v

v = 30 m/s

tension in the string = m * v^2/L

tension in the string = 0.20 * 30^2/0.50

tension in the string = 360 N

the tension in the string is 360 N